OpenStudy (anonymous):

3x + 2y = 7 y = -3 + 11 a.(5,-4) b.(6,-7) c.(-4, 19/2) d.(6,-3) Solve the system of equations using substitution

geerky42 (geerky42):

y = -3 + 11 Well, -3 + 11 = 11 - 3 = 8 y = 8